A. Freyberger, R. Geng, C. Hotaver, G. Krafft, A. Kimber, R. Legg, F. Marhauser, R. Michaud, R. Nelson, T. Plawski, T. Powers, B. Rimmer, C. Reece, T. Reilly, A. Solopova, and H. Wang.
Agenda and Notes
1. Operations Status and Summary/Review (Freyberger)
Notable event during OPS last week. Net result is gradients in two modules turned back to below onset of FE. Question to be discussed is whether all should get similar treatment? More neutron detectors ordered. Do measurements during the summer?
Mounts/Powers observed many quench faults in 1L25-8 that are not quenches but are ????
1L25 and 1L26 thermal cycled to 30 K in Jan. Modest Help?
1L25-8 becomes Unstable
1L25-8 by-passed not recoverable
gradient loss in 1L25-7
Neutron detectors used to find new 1L25 and 1L26 gradients
20+ MeV energy loss in 1L25
8 MeV energy loss in 1L26
Warm region vacuum almost back to nominal
Scaled back to 1000 in the NLINAC, 1020 in the SLINAC
New neutron detectors placed below both 1L25 and 1L26 Tune gradient to no neutrons detected.
2. Cryomodule Shuffle ????
3. Plasma Cleaning (Powers)
Leadership of quenches. Offline discussion.
C100 Refurbishment Project for the future?